Shift report 20030322 00:00-08:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Sat Mar 22 2003 - 07:51:04 EST

  • Next message: Apache: "Shift report 20030322 08:00-16:00"
    ------------------------------------------------------------
    Date: Sat Mar 22 07:51:03 2003, Shift: 00:00-08:00
    Supervisor: sjs, with: 
    Log book#14 pages 71 - 71
    
    Shift Summary:
    ------------------------------------------------------------
    1. Usable beam obtained at 3:00.  Continuous since then.
    2. runs 8601-8603 finishes 12 deg 1/4A runs.
       runs 8581-8583 are for shift checklist
       run 8604 started: D1 and D2 at 1/8A, D3 and D4 off
    
       Updated General Shift Instructions.  In particular, 
       added instructions on how to recover from the problems
       that I had at the beginning of the shift.
    
    3. First found TOFW and some CALFB voltages off during shift
       check. Required rebooting mainframe (and starting it in
       BRAVO).
    
       Then found events not incrementing in DAQ.  Solution was
       to reset TPM2.
    
       Hiro was called to help with both of these problems.  
       Hopefully the revised Problem FAQ will allow him to
       get some sleep if the same problems occur in the future.
    ------------------------------------------------------------
    


    This archive was generated by hypermail 2.1.5 : Sat Mar 22 2003 - 07:55:35 EST